OpenStudy (bunnyzrockgirl):

ANYONE please help me! 17th term: a1 = 2.3; d = -2.3 and 37th term: a1 = -3; d = 2.8

OpenStudy (bunnyzrockgirl):

@Compassionate

OpenStudy (compassionate):

Sorry, not familiar with this type of math. I'll tag a few people. @mathmath333 , @Mathmale , @robtobey , @dan815 , @inkyvoyd

OpenStudy (bunnyzrockgirl):

Thank you!

OpenStudy (bunnyzrockgirl):

@hhelpplzzzz

OpenStudy (anonymous):

\(a_1=2.3\) \(a_2=a_1+d\) \(a_3=a_2+d=a_1+2d\) So I think \(a_n=a_1+(n-1)d\)